OpenStudy (anonymous):

Find the sum using the partial sums. as n approaches infinity.. a_n=1/(n+1) - 1/(n+2)

OpenStudy (dumbcow):

it wants the sum though...however that means it will converge

OpenStudy (anonymous):

yes you are correct didn't read question properly

OpenStudy (anonymous):

does he want this?: \[\sum_{n = 0}^{\infty}a_n\] ? its a telescoping sum.

OpenStudy (dumbcow):

looks like it im rusty at these..so i'll let someone else take over

OpenStudy (anonymous):

I believe the first one converges to 1 then the second converges to 1/2. Therefore 1-1/2= 1/2

OpenStudy (anonymous):

I used the integral test

OpenStudy (anonymous):

If i write out the first n terms i would get: \[(\frac{1}{1}-\frac{1}{2})+(\frac{1}{2}-\frac{1}{3})+(\frac{1}{3}-\frac{1}{4})+\cdots +(\frac{1}{n+1}-\frac{1}{n+2})\] all the terms in the middle will cancel out, leaving: \[1-\frac{1}{n+2}\] So basically you want: \[\lim_{n\rightarrow \infty}1-\frac{1}{n+2} = 1\]

OpenStudy (anonymous):

Wait are you sure the answer is 1 joemath? i dubbed check my work with wolfram and it also said 1/2

OpenStudy (anonymous):

very sure. the integral test doesnt tell you the value of S_n, it just lets you know if it converges or diverges. Since you got an answer using the integral test, that lets you know it converges, but not to what.

OpenStudy (dumbcow):

it might be that you are starting at n=0 instead of n=1

OpenStudy (anonymous):

...or that LOLOLOL

OpenStudy (anonymous):

yup thats it dumb cow

OpenStudy (anonymous):

i did 1 to infinity

OpenStudy (anonymous):

5 medals for you.

OpenStudy (anonymous):

i did 0 to infinity.

OpenStudy (anonymous):

I'm still unsure which one to use though

OpenStudy (anonymous):

was i right in assuming 1 to infinity or is it 0 to infinity

OpenStudy (anonymous):

it wasnt specified. so we are both right unless the poster says otherwise.

OpenStudy (anonymous):

ahhh ok then

OpenStudy (anonymous):

its not like either of us are wrong lol.

OpenStudy (anonymous):

ha ha no i didn't no if these question implicitly are 0 to infinity and i just forgot

OpenStudy (anonymous):

i really depends on the function of n. if it had been something with 1/n in it, i would have started at 1, since n = 0 isnt defined.

OpenStudy (anonymous):

it*

OpenStudy (anonymous):

right good point.